OpenStudy (anonymous):

If MN is the diameter of a circle with M(-3, -4) and N(-7, -14), find the center of the circle

OpenStudy (anonymous):

I'm pretty confused on this question.

OpenStudy (anonymous):

radius=diameter/2; use distance formula \[r=\frac{1}{2}\sqrt{(-14-(-4))^2+(-7-(-3))^2}=.5*\sqrt{10^2+4^2}=.5* \sqrt{116}\]

OpenStudy (anonymous):

since 116=4*29\[=\frac{1}{2}*2*\sqrt{29}=\sqrt{29}\]

OpenStudy (anonymous):

thats not one of the possible answers. the possiblity's are a. -2, -5 b. -4,-10 c. -5,-9 and d. -10, -18

OpenStudy (anonymous):

If the segment MN is the diamter of the circle then we find the distance from M to N and divide by 2

OpenStudy (anonymous):

Sorry, I miss-read the question...i found radius

OpenStudy (anonymous):

The center will be the midpoint of MN \[x=\frac{-3-7}{2}=-5\]and\[y=\frac{-4-14}{2}=-9\]are the coordinates of the center, so it is answer C

OpenStudy (anonymous):

thank you very much :)